Random variable of size n=2 are drawn from a population consisting the numbers 8,10, 12,14,and 16. Construct a sampling distribution of the sample mean to answer the questions.
1
Expert's answer
2021-10-15T09:28:53-0400

I.


"mean=\mu=\dfrac{8+10+12+14+16}{5}=12"

"variance=\sigma^2=\dfrac{1}{5}((8-12)^2+(10-12)^2"

"+(12-12)^2+(14-12)^2+(16-12)^2)=8"

"\sigma=\sqrt{\sigma^2}=\sqrt{8}=2\sqrt{2}"

II. There are "\dbinom{5}{2}=10" samples of size two which can be drawn without replacement: 


"\begin{matrix}
 Sample & Sample\ mean \\
 (8,10) & 9 \\
 (8,12) & 10 \\
 (8,14) & 11 \\
 (8,16) & 12 \\
 (10,12) & 11 \\
 (10,14) & 12 \\
 (10,16) & 13\\
 (12,14) & 13 \\
 (12,16) & 14 \\
 (14,16) & 15 \\
 
\end{matrix}"

III.


"\begin{matrix}
 \bar{X} & P(\bar{X}) \\
 9 & 0.1 \\
 10 & 0.1 \\
 11 & 0.2 \\
 12 & 0.2 \\
 13 & 0.2 \\
 14 & 0.1\\
 15 & 0.1\\

\end{matrix}"

IV.


"\mu_{\bar{X}}=9(0.1)+10(0.1)+11(0.2)+12(0.2)"

"+13(0.2)+14(0.1)+15(0.1)=12"


"\sum_i\bar{X}_i^2P(\bar{X_i})=9^2(0.1)+10^2(0.1)+11^2(0.2)"

"+12^2(0.2)+13^2(0.2)+14^2(0.1)+15^2(0.1)=147"

"\sigma_{\bar{X}}^2=\sum_i\bar{X}_i^2P(\bar{X_i})-\mu_{\bar{X}}^2=147-12^2=3"

"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{3}"

"\mu_{\bar{X}}=12, \sigma_{\bar{X}}=\sqrt{3}"

V.

The mean "\mu_{\bar{X}}" and standard deviation "\sigma_{\bar{X}}" of the sample mean "\bar{X}" satisfy


"\mu_{\bar{X}}=12=\mu,"

"\sigma_{\bar{X}}=\sqrt{3}=\dfrac{2\sqrt{2}}{\sqrt{2}}\sqrt{\dfrac{5-2}{5-1}}=\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}"